OpenStudy (anonymous):

If the radius of the base is 9 units and the height is 11 units, what is the volume of the cylinder shown above?

OpenStudy (anonymous):

area of the base = 2*pi*r area of the cylinder = area of the base * height of the cylinder

OpenStudy (anonymous):

post the pic please

OpenStudy (anonymous):

sorry im an idiot, 2*pi*r is circumference, i was doing the area. pi*r^2 is the area of the base

OpenStudy (anonymous):

so you want 2*pi*r*height = pi*891

OpenStudy (anonymous):

B is the answer